# VisualCoder: Guiding Large Language Models in Code Execution with Fine-grained Multimodal Chain-of-Thought Reasoning

### v2

@@ Abstract
Predicting program behavior and reasoning about code execution remain significant challenges in software engineering, particularly for large language models (LLMs) designed for code analysis. While these models excel at understanding static syntax, they often struggle with dynamic reasoning tasks. We introduce VisualCoder, a simple yet effective approach that enhances code reasoning by integrating multimodal Chain-of-Thought (CoT) reasoning with a visual Control Flow Graph (CFG). By aligning code snippets with their corresponding CFGs, VisualCoder provides deeper insights into execution flow, enabling more accurate predictions of code behavior. Our experiments demonstrate that augmenting LLMs with visual CFGs significantly outperforms text-based CFG descriptions in code reasoning tasks. We address challenges in multimodal CoT integration through a reference mechanism, ensuring consistency between code and its execution path, thereby improving performance in program behavior prediction, error detection, and output generation.
@@
Cuong Chi Le1, Hoang-Chau Truong-Vinh1††thanks: Co-first author., Huy Nhat Phan1, Dung D. Le2, Tien N. Nguyen3, Nghi D. Q. Bui1††thanks: Corresponding author. 1FPT Software AI Center, Viet Nam {cuonglc4, chautvh, huypn16, nghibdq}@fpt.com 2VinUniversity, Vietnam dung.ld@vinuni.edu.vn 3University of Texas at Dallas, USA tien.n.nguyen@utdallas.edu
@@ 1 Introduction
Recent advancements in Code Large Language Models (LLMs) (Hui et al., 2024; Rozière et al., 2024; Wang et al., 2023; Bui et al., 2021; Nijkamp et al., 2023; Lozhkov et al., 2024; Stallone et al., 2024; To et al., 2023) have pushed the boundaries of complex reasoning tasks, extending to the domains that require an understanding of code and its intricacies. There are diverse approaches aimed at enhancing LLMs’ ability. LLMs, while excellent at capturing static patterns and syntax from large code corpora, primarily rely on learned associations rather than direct interaction with the program’s execution environment. They struggle with tasks involving dynamic behaviors of programs, such as predicting execution traces, variable values, or runtime errors, because these tasks require precise understanding of runtime context and program states change during execution. They do not inherently simulate code execution, which is necessary for understanding how variables’ values vary along program’s execution flow. Moreover, LLMs lack the ability to track mutable state or anticipate runtime conditions, leading to difficulties in predicting dynamic behaviors that depends on context-sensitive execution paths.
@@
Recent work has been proposed to enhance the capability of the models in understanding code execution by incorporating Control Flow Graph (CFG) in their reasoning step (Le et al., 2024). It demonstrates that incorporating CFG of given code can significantly improve performance on the code coverage prediction task. However, it utilizes CFGs through graph neural networks rather than directly integrating them into LLM-based reasoning. Despite these advances, most existing work focuses on a single-modality input (i.e., plain code) and has yet to explore the potential of multimodal approaches for code execution reasoning. While code can be read in a linear fashion, understanding its full behavior requires focusing on the non-linear structure of its execution, something that is often visualized more clearly through a CFG.
@@
In recent years, Vision-Language Large Models (VLLMs) (OpenAI et al., 2024; Chen et al., 2023; Liu et al., 2024), have made significant progress, showing their potential across a wide range of tasks that involve both visual and textual inputs. These models, which integrate information from multiple modalities, have been successfully applied to tasks like image captioning, visual question answering, and multimodal retrieval. Recent advancements in multimodal LLMs, such as Flamingo Alayrac et al. (2024), CLIP Radford et al. (2021), and BLIP-2 Li et al. (2023), highlight the benefits of combining visual and textual inputs for enhanced reasoning. Models like LLaVA Liu et al. (2023) and MiniGPT-4 Zhu et al. (2024) show improved performance in multimodal tasks by integrating both visual and textual inputs. Studies have shown that combining visual representations with text significantly improves reasoning, especially in tasks involving complex structures Wei et al. (2024b).
@@
In this work, we propose enhancing the code execution reasoning of LLMs by leveraging multimodal reasoning, combining plain code with visual representations of the corresponding CFG. In our preliminary experiments, simply presenting the plain code alongside textual or visual representations of the CFG has poor performance for code execution-related tasks (Section 5). Recent work by (Zhang et al., 2023) focuses on improving multimodal reasoning in LLMs using the prominent Chain-of-Thought prompting (Wei et al., 2024a) with two separate steps: rationale generation and reasoning to produce answers. However, when applied to our multimodal setup of plain code and CFG, that approach suffers from cascading errors, where inaccuracies in rationale generation negatively impact the reasoning and final answers.
@@
We introduce VisualCoder, a simple yet effective Reference CoT prompting technique that explicitly links individual lines of code to their corresponding visual elements in the CFG. By making these detailed references, our approach encourages the model to focus on specific connections between the code and its execution flow during multimodal reasoning process. This method is expected to improve the LLM’s performance by guiding it to reason more effectively and grounding its reasoning process with more intuitive and informative representation of code graph via imaging, utilizing both the code structure and its execution dynamics.
@@ 2.2 Reasoning about Program Execution
Research into program execution reasoning has advanced through various approaches. They use execution states from constructed programs Chen et al. (2021); Ni et al. (2024); Shin et al. (2018) or predict intermediate subgoals to improve search strategies in sequence-to-sequence models Shi et al. (2024). Another approach trains neural networks to simulate execution, acting as learned interpreters Bieber et al. (2020); nye2021show, often relying on specialized architectures to model flows and data dependencies. Other works like Scratchpad and Self-Debugging explored LLM-generated reasoning chains, while NExT (Ni et al., 2024) uses runtime traces for task-specific rationales.
@@ 3 Motivation
Recent advancements in Large Language Models (LLMs) have shown promise in tasks like code execution prediction, especially with Chain-of-Thought (CoT) reasoning Dhulipala et al. (2024). However, LLMs still struggle with understanding complex execution flows, such as iterations and conditions. Our results in Table 1 (see details later) demonstrate that incorporating Control Flow Graphs (CFG) with source code significantly boosts performance. CFG images provide a visual structure of execution flow, capturing key control structures like branches and loops. This helps LLMs better understand non-linear execution paths, improving program behavior reasoning.
Choosing the right data representation for CFGs is crucial for helping LLMs understand code execution. To support the use of visual representations, we conducted an experiment comparing the effectiveness of textual vs. visual CFGs. As highlighted in Table 2 on our experimental results, the models that utilized visual CFG images consistently outperformed those relying on text-based CFG representation. Our results demonstrate that when models are exposed to CFG images rather than text-based descriptions, their reasoning and prediction accuracy improves substantially. Since text-based representations only provide a linear and sequential description of control flow in textual format, they often fall short in capturing the structural complexity of code execution which requires forward-backward reasoning continuously. In contrast, the visual modality provides an additional layer of information, allowing the model to better comprehend non-linear code flows, such as loops and branches, which are harder to grasp through sequential textual descriptions alone. This result is also consistent with the one in Wei et al. (2024b), which emphasizes that incorporating visual representations significantly enhances the reasoning capabilities of multimodal LLMs. Importantly, this result motivates us on the adopting of visual representations that requires deep, non-linear flow of execution reasoning.
@@
Let us use an example for illustration. As shown in Figure 1, the CFG + CoT approach fails to capture the critical point in reasoning. As with this approach (see red section), the model incorrectly identifies the termination point within the else block (G += 1), missing the fact that this branch is unreachable. Since X is always even, the else block will never execute.
We hypothesize that the key issue is the model’s inability to align the code with its corresponding CFG image during reasoning. Without proper alignment with the CFG, the model misinterprets this unreachable path as a valid termination point, focusing on an irrelevant error. Therefore, we guide the model to refer to each line of code to the corresponding element in the CFG as shown in Figure 1 (highlighted in yellow). Let us call it CFG + CoT + Reference approach, which correctly identifies the unreachable node and termination point. Our results (Section 5) also show that the two-stage multimodal-CoT approach in Zhang et al. (2023) is also insufficient for complex coding tasks that involve intricate execution flows.
As illustrated in Figure 1, the CFG + CoT + Reference approach (green section) allows the LLM to correctly identify the critical point: the unreachable nature of the else branch. By explicitly referencing the CFG during reasoning, the model avoids errors in unreachable branches and focuses on the actual critical error—the float N being used in the range() function. This reference mechanism helps the model maintain proper alignment between the visual CFG and the code, leading to more accurate reasoning on code execution at runtime.
@@ 4 VisualCoder: Reference Mechanism
We propose a method that combines Control Flow Graphs (CFG) with Chain-of-Thought (CoT) reasoning, augmented by a Reference Mechanism, to facilitate enhanced code execution reasoning. This approach enables step-by-step evaluation of the code while also cross-referencing control flow points, thereby improving error detection and identifying unreachable or erroneous code paths.
Let the Python code snippet be represented as a sequence of lines of code:
@@ 4.1 Chain-of-Thought Reasoning (CoT)
Chain-of-Thought reasoning is implemented by analyzing each instruction on while considering its logical dependencies. We define the reasoning process as a recursive function:
where is a function that takes the current line of code and the context, iterating through each step of the code while considering the nodes in the CFG.
@@ 4.2 Reference Mechanism
The Reference Mechanism augments the CoT reasoning by mapping each line of code to its respective CFG node. This mapping is expressed as: , where is represented by node in the CFG. The model now references during reasoning to ensure consistency between the plain code and control flow. This alignment helps the model analyze each line within the control flow, improving its understanding of execution paths, transitions, and dependencies across statements and blocks, rather than treating lines in isolation.
@@ 4.4 CFG + CoT + Reference
Where is the mapped node in the CFG corresponding to the current line . By analyzing and referencing the corresponding CFG block for every line of code, the model can maintain consistency between the control flow and the sequential code.
@@ 4.5 VisualCoder
There are several ways to achieve the behavior outlined in the CFG + CoT + Reference process, such as fine-tuning, one-shot or few-shot prompting, and more. In our current implementation, we propose a straightforward yet effective approach that can be integrated into any Chain-of-Thought framework without the need for fine-tuning. By introducing a simple instruction, as shown in Figure 1 (green line in the prompt), we expect to guide Vision Language Models to follow the formulation described in Equation 5. This approach ensures that the model focuses its reasoning on the relevant CFG node for each line of code, thereby improving its alignment with the control flow. The experimental results in Section 5, along with the qualitative analysis in Section 6 and attention heat map in 7, demonstrate the effectiveness of our method in enhancing code execution reasoning.
5 Experiments
5.1 Better Code Execution Understanding with Control Flow Graph
In this experiment, we aimed to show that providing the LLM with a CFG improves its understanding of code execution. Using the CRUXEval benchmark Gu et al. (2024), we tested models on predicting execution outputs. We compared the performance of three state-of-the-art VLM models—Claude Sonnet 3.5 Anthropic (2024), Gemini-1.5-Flash Reid et al. (2024), and InterVL2-8B Chen et al. (2023)—in two settings: 1) plain code, and 2) plain code with its CFG image. The task involved both output prediction (predicting the execution’s result) and input prediction (predicting the inputs leading to a specific output).
For direct comparison with prior work, we used the same prompt format from the original CRUXEval paper Gu et al. (2024). The prompt provided the code and, when applicable, a visual CFG, guiding step-by-step reasoning. Performance was measured using the Accuracy@1 metric, indicating if the models’ first predictions were correct.
| Task | Settings | Models | Accuracy@1 |
| Plain code | Claude Sonnet 3.5 | 79.6 | |
| Plain code + CFG image | Claude-3.5-Sonnet | 82.3 | |
| Plain code | Gemini-1.5-Flash | 68.5 | |
| Output Prediction | Plain code + CFG image | Gemini-1.5-Flash | 70.0 |
@@ 5.1 Better Code Execution Understanding with Control Flow Graph
| Plain code | Claude Sonnet 3.5 | 75.2 | |
| Plain code + CFG image | Claude Sonnet 3.5 | 84.0 | |
| Plain code | Gemini-1.5-Flash | 58.4 | |
| Input Prediction | Plain code + CFG image | Gemini-1.5-Flash | 68.4 |
@@
The results in Table 1 show that incorporating a CFG image improves model accuracy in two settings. This improvement is consistent across models, showing that CFG enhances the LLMs’ ability to reason about execution flow and predict program behaviors more accurately. This result is consistent with the one reported by (Le et al., 2024) in which incorporating CFG of given code can improve performance on code coverage prediction.

### v3

@@ Abstract
Predicting program behavior and reasoning about code execution remain significant challenges in software engineering, particularly for large language models (LLMs) designed for code analysis. While these models excel at understanding static syntax, they often struggle with dynamic reasoning tasks. We introduce VisualCoder, a simple yet effective approach that enhances code reasoning by integrating multimodal Chain-of-Thought (CoT) reasoning with a visual Control Flow Graph (CFG). By aligning code snippets with their corresponding CFGs, VisualCoder provides deeper insights into execution flows. We address challenges in multimodal CoT integration through a reference mechanism, ensuring consistency between code and its execution path, thereby improving performance in program behavior prediction, error detection, and output generation. Our implementations are available at https://github.com/FSoft-AI4Code/VisualCoder.
@@
Cuong Chi Le1, Hoang-Chau Truong-Vinh1, Huy Nhat Phan1, Dung D. Le2, Tien N. Nguyen3, Nghi D. Q. Bui1††thanks: Corresponding author. 1FPT Software AI Center, Viet Nam {cuonglc4, chautvh, huypn16, nghibdq}@fpt.com 2VinUniversity, Vietnam dung.ld@vinuni.edu.vn 3University of Texas at Dallas, USA tien.n.nguyen@utdallas.edu
@@ 1 Introduction
Recent advances in Code-related Large Language Models (LLMs) (Hui et al., 2024; Rozière et al., 2024; Wang et al., 2023; Bui et al., 2021; Nijkamp et al., 2023; Lozhkov et al., 2024; Stallone et al., 2024; To et al., 2023; Guo et al., 2024; Wei et al., 2024c; Manh et al., 2024; Huang et al., 2024; Muennighoff et al., 2023) have pushed the boundaries of complex reasoning tasks, extending to the domains that require an understanding of code and its intricacies. There are diverse approaches aimed at enhancing LLMs’ ability. LLMs, while excellent at capturing static patterns and syntax from large code corpora, primarily rely on learned associations rather than direct interaction with the program’s execution environment. They struggle with tasks involving dynamic behaviors of programs, such as predicting execution traces, variable values, or runtime errors, because these tasks require precise understanding of runtime context and program states change during execution. They do not inherently simulate code execution, which is necessary for understanding how variables’ values vary along program’s execution flow. Moreover, LLMs lack the ability to track mutable state or anticipate runtime conditions, leading to difficulties in predicting dynamic behaviors that depends on context-sensitive execution paths.
@@
Recent work has been proposed to enhance the capability of the models in understanding code execution by incorporating Control Flow Graph (CFG) in their reasoning step (Le et al., 2024; Bieber et al., 2020, 2023). It demonstrates that incorporating CFG of given code can significantly improve performance on the code coverage prediction task. However, it utilizes CFGs through graph neural networks rather than directly integrating them into LLM-based reasoning. Despite these advances, the state-of-the-art approaches focus on a single-modality input (i.e., plain code) and has yet to explore the potential of multimodal frameworks for code execution reasoning. While code can be read in a linear fashion, understanding its behavior requires focusing on the non-linear flow of execution, which is visualized more clearly via a CFG.
@@
In recent years, Vision Language Large Models (VLLMs) (OpenAI et al., 2024; Chen et al., 2024; Liu et al., 2024), have made significant progress, showing their potential across a wide range of tasks that involve both visual and textual inputs. These models, which integrate information from multiple modalities, have been successfully applied to tasks like image captioning, visual question answering, and multimodal retrieval. Recent advances in multimodal LLMs, such as Flamingo Alayrac et al. (2024), CLIP Radford et al. (2021), and BLIP-2 Li et al. (2023a), highlight the benefits of combining visual and textual inputs for enhanced reasoning. Models like LLaVA Liu et al. (2023) and MiniGPT-4 Zhu et al. (2024) show improved performance in multimodal tasks by integrating both visual and textual inputs. Studies have shown that combining visual representations with text significantly improves reasoning, especially in tasks involving complex structures Wei et al. (2024b).
@@
In this work, we propose enhancing the program execution reasoning of LLMs by leveraging multimodal reasoning, combining plain code with visual representations of the corresponding CFG. In our experiments, simply presenting the plain code alongside textual or visual representations of the CFG has poor performance for program execution-related tasks (Section 5). Recent work by (Zhang et al., 2023) focuses on improving multimodal reasoning in LLMs using the prominent Chain-of-Thought prompting (Wei et al., 2024a) with two separate steps: rationale generation and reasoning to produce answers. However, when applied to our multimodal setup of plain code and CFG, that approach suffers from cascading errors, where inaccuracies in rationale generation negatively impact the reasoning and final answers.
@@
We introduce VisualCoder, a simple yet effective Reference CoT prompting technique that explicitly links individual lines of code to their corresponding visual elements in the CFG. By making these detailed references, our approach encourages the model to focus on specific connections between the code and its execution flow during multimodal reasoning process. This technique is expected to improve the LLM’s performance by guiding it to reason more effectively and grounding its reasoning process with more intuitive and informative representation of code graph via imaging, utilizing both the code structure and its execution dynamics.
@@ 2 Related Work
2.1 Code Large Language Models
Large Language Models (LLMs) have been widely applied to various code-related tasks, including code understanding, reasoning, and analysis (Chen et al., 2021a; Li et al., 2023b; Jiang et al., 2024; Touvron et al., 2023; Rozière et al., 2024; Xu et al., 2022; Allal et al., 2023; Nijkamp et al., 2022; Phan et al., 2024a; To et al., 2023; Manh et al., 2024; Phan et al., 2024b). Early benchmarks Yin et al. (2018); Iyer et al. (2018); Nguyen et al. (2023); Chen et al. (2021a); Austin et al. (2021); Hendrycks et al. (2021) primarily assessed model performance using match-based similarity metrics, which fail to capture deeper reasoning and functional correctness Chen et al. (2021a). Some benchmarks emphasize domain diversity Yin et al. (2018); Iyer et al. (2018); Nguyen et al. (2023), while others, such as HumanEval Chen et al. (2021a), MBPP Austin et al. (2021), and APPS Hendrycks et al. (2021), focus on specific tasks like function completion or competitive programming. More recent efforts have sought to expand the scope of evaluation, such as ExeDS Huang et al. (2022), which targets data science workflows, and ODEX Wang et al. (2022), an open-domain evaluation suite. However, these benchmarks primarily assess static code properties and standalone function reasoning, with limited emphasis on execution flow analysis and dynamic behavior prediction—critical aspects for improving LLMs’ ability to reason about code execution.
@@ 2.3 Reasoning about Program Execution
Research into program execution reasoning has advanced through various approaches. They use execution states from constructed programs Chen et al. (2021b); Ni et al. (2024); Shin et al. (2018) or predict intermediate subgoals to improve search strategies in sequence-to-sequence models Shi et al. (2024). Another approach trains neural networks to simulate execution, acting as learned interpreters Bieber et al. (2020, 2023); Le et al. (2024), often relying on specialized architectures to model flows and dependencies. Other works like Scratchpad and Self-Debugging explored LLM-generated reasoning chains, while NExT (Ni et al., 2024) uses runtime traces for task-specific rationales.
@@ 3 Motivation
Recent advances in Large Language Models (LLMs) have shown promise in tasks like program execution prediction, especially with Chain-of-Thought (CoT) reasoning Dhulipala et al. (2024). However, LLMs still struggle with understanding complex execution flows, such as iterations and conditions. Our results in Table 1 (see details later) demonstrate that incorporating Control Flow Graphs (CFG) with source code significantly boosts performance. CFG images provide a visual structure of execution flow, capturing key control structures like branches and loops. This helps LLMs better understand non-linear execution paths, improving program behavior reasoning.
Choosing the right data representation for CFGs is crucial for helping LLMs understand code execution. To support the use of visual representations, we conducted an experiment comparing the effectiveness of textual vs. visual CFGs. As highlighted in Table 2 on our experimental results, the models that utilized visual CFG images consistently outperformed those relying on text-based CFG representation. Our results demonstrate that when models are exposed to CFG images rather than text-based descriptions, their reasoning and prediction accuracy improves substantially.
Since text-based representations only provide a linear and sequential description of control flow in textual format, they often fall short in capturing the structural complexity of code execution which requires forward-backward reasoning continuously. In contrast, the visual modality provides an additional layer of information, allowing the model to better comprehend non-linear code flows, such as loops and branches, which are harder to grasp through sequential textual descriptions alone. This result is also consistent with the research by Wei et al. Wei et al. (2024b), which emphasizes that incorporating visual representations significantly enhances the reasoning capabilities of multimodal LLMs. Importantly, this result motivates us on the adopting of visual representations that requires deep, non-linear flow of execution reasoning.
@@
Let us use an example for illustration. As shown in Figure 1, the CFG + CoT approach fails to capture the critical point in reasoning. As with this approach (see red section), the model incorrectly identifies the termination point within the else block (G += 1), missing the fact that this branch is unreachable. Since X is always even, the else block will never be executed.
We hypothesize that the key issue is the model’s inability to align the code with its corresponding CFG image during reasoning. Without proper alignment with the CFG, the model misinterprets this unreachable path as a valid termination point, focusing on an irrelevant error. Therefore, we guide the model to refer to each line of code to the corresponding element in the CFG as shown in Figure 1 (highlighted in green). Let us call it CFG + CoT + Reference approach, which correctly identifies the unreachable node and termination point. Our results (Section 5) also show that the two-stage multimodal-CoT approach in Zhang et al. (2023) is also insufficient for complex coding tasks that involve intricate execution flows.
As illustrated in Figure 1, the CFG + CoT + Reference approach (green section) allows the LLM to correctly identify the critical point: the unreachable nature of the else branch. By explicitly referencing the CFG during reasoning, the model avoids errors in unreachable branches and focuses on the actual critical error—the float N being used in the range() function. This reference mechanism helps the model maintain proper alignment between the visual CFG and the code, leading to more accurate reasoning on program execution at runtime.
@@ 4 VisualCoder: Reference Mechanism
We propose a method that combines Control Flow Graphs (CFG) with Chain-of-Thought (CoT) reasoning, augmented by a Reference Mechanism, to enhance reasoning on program execution. This approach enables step-by-step evaluation of the code while also cross-referencing control flow points, thereby improving error detection and identifying unreachable or erroneous code paths.
Let the given Python code snippet be represented as a sequence of lines of code:
@@ 4.1 Chain-of-Thought Reasoning (CoT)
Chain-of-Thought reasoning is implemented by analyzing each instruction on while considering its logical dependencies and its corresponding control flow in the CFG. We define the reasoning process as a recursive function:
where is a function that takes the current line of code, its execution context, and its corresponding CFG node .
@@ 4.2 Reference Mechanism
The Reference Mechanism enhances CoT reasoning by mapping each line of code to its corresponding CFG node, expressed as , where is represented by node in the CFG. To establish this mapping, we guide the model to focus on the relevant CFG node while reasoning about each line of code. This is achieved by reinforcing attention on the corresponding node in the CFG image whenever the model processes its associated code line, as demonstrated in Section 7. By aligning each line of code with its node in the control flow representation, this mechanism improves the model’s understanding of execution paths, transitions, and dependencies across statements and blocks, rather than treating lines in isolation.
@@ 4.4 CFG + CoT + Reference
Where is the mapped node in the CFG corresponding to the current line . By analyzing and referencing the corresponding CFG block for every line of code, the model can maintain consistency between the CFG and the source code.
@@ 4.5 VisualCoder
There are several ways to achieve the behavior outlined in the CFG + CoT + Reference process, such as fine-tuning, one-shot or few-shot prompting, and more. In our current implementation, we propose a straightforward yet effective approach that can be integrated into any Chain-of-Thought framework without the need for fine-tuning. By introducing a simple instruction, as shown in Figure 1 (green line in the prompt), we expect to guide Vision Language Models to follow the formulation described in Equation 5. This approach ensures that the model focuses its reasoning on the relevant CFG node for each line of code, thereby improving its alignment with the control flow. The experimental results in Section 5, along with the qualitative analysis in Section 6 and attention heat map in Section 7, demonstrate the effectiveness of our method in enhancing program execution reasoning.
5 Empirical Evaluation
5.1 Better Code Execution Understanding with Control Flow Graph Images
In this experiment, we aimed to show that providing the LLM with CFG images (no references) improves its code execution reasoning. Using the CRUXEval benchmark Gu et al. (2024), we tested models on predicting execution outputs. We compared the accuracies of three state-of-the-art VLM models—Claude 3.5 Sonnet Anthropic (2024), Gemini-1.5-Flash Reid et al. (2024), and InterVL2-8B Chen et al. (2024)—in two settings: 1) plain code, and 2) plain code with its CFG image. The task involved both output prediction (predicting the execution’s result) and input prediction (predicting the inputs leading to a specific output).
For direct comparison with prior work, we used the same prompt format from the original CRUXEval paper Gu et al. (2024). The prompt provided the code and, when applicable, a visual CFG, guiding step-by-step reasoning. Performance was measured using the pass@1 metric, indicating if the models’ first predictions were correct.
| Task | Settings | Models | pass@1 |
| Plain code | Claude 3.5 Sonnet | 79.6 | |
| Plain code + CFG image | Claude 3.5 Sonnet | 82.3 | |
| Plain code | Gemini 1.5 Flash | 68.5 | |
| Output Pred. | Plain code + CFG image | Gemini 1.5 Flash | 70.0 |
@@ 5.1 Better Code Execution Understanding with Control Flow Graph Images
| Plain code | Claude 3.5 Sonnet | 75.2 | |
| Plain code + CFG image | Claude 3.5 Sonnet | 84.0 | |
| Plain code | Gemini 1.5 Flash | 58.4 | |
| Input Pred. | Plain code + CFG image | Gemini 1.5 Flash | 68.4 |
@@
The results in Table 1 show that incorporating a CFG image improves model accuracy in two settings. This improvement is consistent across models, showing that CFG enhances the LLMs’ ability to reason about execution flow and predict program behaviors more accurately. This result is consistent with the one reported by Le et al. (Le et al., 2024) in which incorporating CFG of given code can improve performance on code coverage prediction.